# JEL: C81, O32

### v2

@@
A recent literature contends that the returns to research in medicine are declining. This conclusion is premised on two empirical findings: the quantity of medical research is rising, while improvements in health outcomes are stagnant. We argue that the first finding is an artifact of conceptual and empirical problems in the measurement of research. We propose a method for annotating unstructured text based on fine-tuning a large language model. We apply this method to construct a census of published clinical trials. We argue that clinical trials function as a natural proxy for productivity-relevant research in medicine. In our data, the quantity, quality, and composition of clinical trials have been stable since 2010. We find that measured increases in the total quantity of medical research are driven by shifts in the composition of scientific publications that do not correspond to real increases in the quantity of productivity-relevant research. We document large increases in the growth of non-trial publications, driven by publications from China and publications that review existing literature. This increase in quantity is coincident with a decline in various measures of average publication quality. Taken together, we find no support for declining medical research productivity over this period.
Keywords: Medical Innovation, Productivity, Clinical Trials, Large Language Models, Generative AI
@@ 1. Introduction
Two stylized facts support the argument that the returns to medical research are diminishing. [^1] Goldin et al., (2024) survey evidence related to declining productivity across the economy, including evidence of diminishing returns to research investment in sectors beyond medicine. The quantity of resources directed toward medical research, measured in dollars and scientific publications, appears to be increasing (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020; Park et al.,, 2023). [^2] See Park et al., (2023), Bloom et al., (2020), and Scannell et al., (2012) for case studies examining the quantity of investment in science, medicine, and drug development, respectively. Bloom et al., (2020) collect case studies from other research-intensive fields. See, also, Chu and Evans, (2021), who document large increases in the quantity of papers in certain scientific fields and examine associated patterns in citation networks. Meanwhile, improvements in human health (e.g., increases in life expectancy) continue at the same steady pace (Mehta et al.,, 2020; Ho and Hendi,, 2018). If more resources are necessary to yield the same rates of technological progress, simple models of “idea production” imply that research productivity—the efficiency with which research is converted into new and useful knowledge—must be falling (Jones,, 1995; Bloom et al.,, 2020). [^3] In some models, these patterns can be reconciled without diminishing returns to research. For instance, Young, (1998) investigates the idea that investment in new varieties of existing products can yield the same trends.
These facts, and their forecast of technological stagnation in a sector closely linked to health and welfare (Cutler and McClellan,, 2001; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007), have influenced prominent policy debates. [^4] These facts were most prominently documented in Bloom et al., (2020). Concerns about technological stagnation in modern policy discourse predate Bloom et al., (2020) and respond in part to Cowen, (2011) and Gordon, (2017). As we discuss below, the prospect of diminishing returns to scientific research has long been a subject of concern. Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S. Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that science and medicine are in a state of decline. [^5] This concern has been explored extensively in the popular press. Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), among others, interpret these facts in the popular press as a “warning sign” for the future of technological progress (Piper,, 2023).
In this paper, we find that one of these stylized facts—the increase in the quantity of resources directed toward medical research—is an artifact of conceptual and empirical problems in the measurement of medical research. An ideal, and infeasible, measure of the quantity of research captures the amount of “effort” directed toward increasing the stock of useful (medical) knowledge. Informative proxies for this measure have two characteristics. First, as researchers differ in skill and research designs differ in quality, the proxy must account for heterogeneity in the likelihood that research projects yield “productivity-relevant” information. Second, units of the proxy must be comparable over time. If the quantity of research embedded in, say, a patent or paper decreases with time, increases in the quantity of patents or papers provide very little insight into research productivity.
We find that existing measures of medical research fall short on both accounts. The primary contribution of this paper is the construction of a measure of medical research—a census of published clinical trials—that addresses both issues and fills this gap. We show that standard data construction tools classify research imprecisely, and that this imprecision yields inaccurate conclusions about trends in productivity. Thus, we introduce a method for annotating unstructured text at scale. When applied to a corpus of scientific text, this method allows us to identify and classify heterogeneous forms of research, relaxing a long-standing data constraint in the economics of innovation.
We fine-tune an open-source large language model and devise a workflow that allows us to identify clinical trial records from the unstructured text of scientific publications with an accuracy comparable to that of a human labeler. In our data, the quantity, quality, and composition of clinical trials have been stable since 2010. We find evidence that measured increases in the quantity of scientific publications capture shifts in scientific incentives that are not indicative of real increases in the quantity of (productivity-relevant) research. Taken together, we find no support for claims of declining medical research productivity over this period.
We begin by arguing that clinical trials are a natural proxy for the research object of interest in this setting. Trials are discrete, regulated experiments that generate useful statistical information about new and existing technologies (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024). These experiments are designed in consultation with regulators after sufficiently robust preliminary evidence has been produced but before health technologies can be marketed. Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015). Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications. [^6] In the 1990s and 2000s, medical experts created a checklist that provides authors of papers reporting the results of clinical trials with a standardized structure and list of elements for publication. The most recent Consolidated Standards of Reporting Trials (CONSORT) Statement was published in 2010 and is now the reporting standard for most scientific journals that publish the results of trials. See Laine et al., (2007); Schulz et al., (2010) for details. To our knowledge, this type of standardization is unique to this form of science.
No existing dataset allows us to track quantities of clinical trials over time. Although dollar-denominated expenditures are an obvious metric, available records of investment in trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (e.g., DeVito et al.,, 2020). Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are biased by these same noncompliance patterns.
We construct a new census of clinical trials indexed in scientific publications. We collect all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects. This definition is intentionally restrictive and aims to include only those records that clearly report trials. Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeler.
@@
Manually identifying clinical trial records, at this scale, is infeasible. [^7] A small number of papers, including Oostrom, (2024), examine medical investment using small samples of clinical trial records that are constructed by hand. Standard methods perform poorly. In particular, we show that documented increases in the quantity of clinical trials during this time period, as reported by Bloom et al., (2020), are driven by a categorization error caused by keyword-based searches. [^8] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews). We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time. In our period of interest, this generates a spurious increase in the measured quantity of trials. Machine learning methods used in the health services literature similarly lack the precision necessary for our purposes (Thomas et al.,, 2021). Instead, we leverage recent advances in generative AI, which allow for data extraction from unstructured text at scale. We devise a workflow for extracting precisely defined objects from a large corpus of text, and train a bespoke large language model that is specifically optimized for this task.
@@
We proceed in stages. We devise an interface to iteratively revise model prompts. With the highest-performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023). These noisy labels are used as training data to fine-tune a set of open-source models. An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
@@
These data allow us to document two sets of facts about the quantity of medical research over time. First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest. We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data. On these dimensions, we find considerable heterogeneity. Roughly half of clinical trial publications in our sample are never cited by a leading medical journal. Nearly 15 percent are never cited by any other scientific publication. However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging. Taken together, we interpret these patterns as evidence that “productivity-relevant” research in medicine is essentially constant in quantity and stable in composition over our time period. This contradicts existing findings, at least for our period of interest, and calls into question the conclusion that the productivity of medical research is falling.
Second, our data allow us to examine the source of measured increases in the quantity of scientific publications (Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021). We contrast clinical trial production with trends in the publication of other forms of medical research. While the quantity of clinical trials is constant across our period of interest, the quantity of other medical research—defined, in our preferred version, as the set of papers citing clinical trials—has increased by nearly a factor of two. [^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews. In doing so, we confirm that our focus on clinical trials is appropriate. These increases in quantity have coincided with shifts in the quality and composition of these other types of research. We highlight three patterns. First, the large increase in the total number of medical papers in our sample is largely driven by researchers in China. Second, this rise in quantity coincides with a decrease (on various measures) in the average quality of publications. Third, there is a substantial increase in the quantity of meta-analyses and literature reviews during our period of interest. [^10] These patterns are consistent with the geometric increase in the quantity of meta-analyses produced by Chinese researchers, documented in Ioannidis et al., (2013). We interpret these patterns as evidence that—for this category of other medical research—the quantity of productivity-relevant research “embedded” in a single document may be declining over time. Quantities of publications, then, provide essentially no insight on the question of research productivity. [^11] Park et al., (2023) examine the “disruptiveness” of scientific research, using data on the universe of publications across six decades. Their findings are similar to ours, in the sense that they document that the absolute counts of highly disruptive papers remain constant over time, while average “disruptiveness” is declining. They interpret this as evidence that science is becoming less disruptive over time. Our paper differs in two important ways. First, we distinguish between types of publications on the basis of the economic value they produce. Second, we document that measured increases in the quantity of publications are disproportionately concentrated in types of publication that focus on diffusion, rather than production, of new information. Pooled averages, then, may be difficult to interpret.
Our findings speak directly to ongoing policy discussions and branches of the macroeconomic literature, which consider the hypothesis that the rate of technological progress may be declining. Nearly all micro-level evidence on rising levels of research investment comes from studies of scientific publications and, in particular, clinical trials. [^12] See Goldin et al., (2024) for a summary. Notable exceptions are studies of agriculture and semiconductors. We show that these findings result from conceptual and empirical challenges in the quantification of innovation. Thus, our work is most closely related to a series of papers—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that view the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014). [^13] Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024). More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
In doing so, we draw on, and contribute to, a long literature on the measurement of invention. Pakes and Griliches, (1980) begin a discussion of inventive activity led by U.S. firms by remarking: “Perhaps the most serious task facing empirical work in the area of ‘technological change’ and ‘invention and innovation’ is the construction and interpretation of measures (indices) of advances in knowledge.” In a more recent survey, Bryan and Williams, (2021) note that, although much progress has been made over half a century of research, construction of such measures remains an open challenge. Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation. An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents. Kogan et al., (2017) and Kelly et al., (2021) provide two recent, novel, alternatives. Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends. While scientific papers are a natural alternative to patents on the input side, existing methods for the use of patent data are not easily applied to publications (Popp,, 2015; Popp et al.,, 2010). [^14] In particular, citations across scientific papers are not analogous to citations across patents. Whereas patent citations are legal objects, which are defined in part by patent examiners and delineate the boundaries of a right enforceable by courts, citations across scientific papers reflect norms and networks. See Hall et al., (2000).

### v3

@@
We develop a method for assigning high-quality labels to unstructured text. The method is based on fine-tuning an open-source language model with data extracted from a proprietary large language model. We apply this method to construct a census of published clinical trials. We revisit a literature that contends that pharmaceutical research productivity is declining, based on measured increases in the quantity of clinical trials, which outpace trends in output. In our data, the quantity and composition of trials are stable since 2010. Previous measurements are an artifact of biases driven by shifts in the composition of other forms of research.
Keywords: Large Language Models, Medical Research Productivity, Clinical Trials
@@
Perhaps the most serious task facing empirical work in the area of “technological change” and “invention and innovation” is the construction and interpretation of measures (indices) of advances in knowledge. — Pakes and Griliches,, 1980
@@ 1. Introduction
Considerable progress in the measurement of innovation has been made in the half-century since the publication of Pakes and Griliches, (1980). This progress has been enabled in large part by the proliferation of large-scale corpora of patents and scientific publications. Yet measurement remains an active constraint for empirical research (Bryan and Williams,, 2021). Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research. Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994). Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022). Correction of this confounding can be viewed as a classification problem: how accurately can unstructured text be mapped to specific quantities of interest? [^1] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001). Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category.
In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget. We develop a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification. The class of methods that we consider has two steps. First, a comparatively small number of labels are extracted from a large, typically proprietary, language model. [^2] By proprietary large language model, we are referring, at present, to models such as OpenAI’s GPT 3.5 and GPT-4. Alternative models include Anthropic’s Claude 3.5 Opus, Google’s Gemini, among many others. Second, these labels are used as training data for a smaller, more efficient, open-source language model. [^3] Here, we are referring to classes of open-source language models such as RoBERTa, based on Google’s BERT model, among others. This process is an instance of model distillation, which is a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024). This procedure allows researchers to capture the benefits of complex, frontier models, while retaining the resource efficiency and transparency of lighter-weight open-source models.
The primary contribution of this paper is the application of this procedure in a context that is emblematic of the econometric challenges faced by studies of innovation: measurement of trends in clinical trial production. We show that standard methods for data classification identify clinical trials in publication data imprecisely. These imprecise measurements indicate, spuriously, that the quantity of clinical trials has increased sharply since 2010. In our data, the quantity, quality, and composition of clinical trials have been stable year-on-year.
When considered alongside recent evidence on the stability of drug development costs during this period (Sertkaya et al.,, 2024), these new measurements of clinical trial production call into question the conclusions of an influential literature that finds diminishing returns to pharmaceutical research (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020). [^4] Goldin et al., (2024) survey evidence related to declining productivity across the economy and note that, in the pharmaceutical sector, evidence for declining productivity primarily relies on measures of total numbers of clinical trials and estimates of R&D expenditures. The conclusions of this literature are based on measures that suggest that the quantity of clinical trials is increasing substantially, in a way unmatched by trends in measures of outputs. In light of our new data, we find that there is little rigorous empirical support for falling pharmaceutical research productivity in the most recent period. This finding is directly relevant to a set of active policy debates, which were initiated, in no small part, in response to concerns that the production of innovative output in this sector is slowing down. These debates include proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S. Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022).
Clinical trials are an especially attractive measure of research in medicine (Bryan and Williams,, 2021). In the pharmaceutical industry, clinical trials are discrete, regulated experiments that generate useful statistical information about new and existing medicines (Chavez-MacGregor and Giordano,, 2016). Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015). Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021). Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see Finkelstein, (2004) and Budish et al., (2015) for two influential examples). Studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)). Recently, and prominently, measures of clinical trial quantities have been used to support arguments that medical and pharmaceutical research productivity are declining over time (Scannell et al.,, 2012; Bloom et al.,, 2020).
Nevertheless, it is remarkably difficult to generate accurate measures of trends in clinical trial production. Administrative databases—such as ClinicalTrials.gov—have experienced large shifts in compliance with reporting regulations over time, which make it difficult to distinguish genuine increases in research effort from greater adherence to reporting requirements (see e.g., DeVito et al.,, 2020). Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are confounded by the same shifts in compliance.
@@
Given these limitations, scientific publication data have served as a natural alternative or complement in recent empirical studies (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024). With unstructured publication text data, the challenge becomes, again, one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial? Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible. [^5] A small number of papers curate select samples of clinical trial data, for specific research questions. For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy. We show that standard methods—based on publication metadata and machine learning tools—perform poorly. In particular, we show that keyword-based searches, including methods employed by Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials. [^6] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews). We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time. Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends.
Instead, we leverage recent advances in generative AI, which enable data extraction from unstructured text at scale. Our objective is to identify all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects. We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler. [^7] Many records in PubMed / MEDLINE have full-texts available for download, via the PubMed Central database. However, these full-texts are subject to copyright restrictions. Roughly one quarter of records in PubMed—$8$ million of the $34$ million total records—have full-texts that are available under licenses that would permit this type of research use (the PubMed Central Open Access Subset, at https://www.ncbi.nlm.nih.gov/pmc/tools/openftlist/). Thus, we restrict our focus to abstracts, for which no such restrictions apply.
@@
We proceed in stages. Given the abstract of a scientific publication, we prompt two proprietary language models, OpenAI’s GPT-3.5 and GPT-4 (Ouyang et al.,, 2022; Achiam et al.,, 2023; Bubeck et al.,, 2023), to label whether the record satisfies our desired sample restrictions. We find considerable variation in performance with seemingly small differences in prompt text. We iteratively revise prompt text until small changes yield no improvement in performance. GPT-4 performs quite well off-the-shelf. The performance of GPT-3.5 is unsuitable for our application. However, use of GPT-4 at scale—to classify millions of records—is computationally and financially impractical. We classify a moderate number of abstracts with the best-performing prompts. The labels extracted from the proprietary language models are used as training data to fine-tune a set of open-source models. Our preferred sample is constructed using an ensemble of fine-tuned models, which matches the performance of the best proprietary model. False positive and false negative rates are below 5 percent. When this model is used to label the full sample of PubMed records, it yields a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
Three technical features of this fine-tuning process are worth highlighting. First, we document a phase-transition in model quality in the number of training labels used. Put differently, we observe a threshold in the number of training labels used in fine-tuning, beyond which the performance of the fine-tuned model increases dramatically. Second, models fine-tuned with labels extracted from GPT 3.5 exceed the performance of GPT 3.5. Third, models fine-tuned with labels extracted from GPT-4 match the performance of GPT-4 itself. That is, we show that researchers can distill the full benefits of large proprietary models into light-weight, resource-efficient models that can be realistically applied to standard data construction tasks.
@@
The resulting data allow us to document two sets of facts about the quantity of medical research over time. First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest. We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data. On these dimensions, we find considerable heterogeneity that is, to our knowledge, new to the literature. Roughly half of clinical trial publications in our sample are never cited by a leading medical journal. Nearly 15 percent are never cited by any other scientific publication. However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging.
Second, our data allow us to shed light on the source of the classification errors that drive the trends documented in Bloom et al., (2020). [^8] These data are relevant outside of this particular context, for studies of medical innovation and for studies in the health services literature that use clinical trial publications to construct prescribing guidelines for physician practice. Identifying trial records in publication data has been a challenge for this health services literature, where existing best practices have both high true and false positive rates. See Thomas et al., (2021). As similar trends in the total quantity of scientific publications have been documented elsewhere, including in Park et al., (2023) and Chu and Evans, (2021), these insights may be more generally applicable to discussions about how large increases in the quantity of publications should be interpreted. We contrast trends in clinical trial production with trends in the production of medical research that is similar—conceptually and in its textual presentation in scientific publications—to clinical trials. In our preferred version, this set of other medical research is defined as the set of papers that cite clinical trials. [^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews. While the quantity of clinical trials is constant over this period, the quantity of this other non-trial research has increased by a factor of two.
Three facts about shifts in the composition and quality of these publications provide insight into the source of measured increases in research quantity. First, the large increase in the total number of non-trial papers in our sample is largely driven by researchers in China. Second, this rise in quantity coincides with a decrease (on available measures) in the average quality of publications. Third, there is a substantial increase in the quantity of meta-analyses and literature reviews. The main conclusion that emerges from this analysis is that errors in the classification of scientific publications can yield spurious findings about trends in research effort.
1.1. Related Literature
We draw on, and contribute to, a long literature on the measurement of invention. Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation. An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents. Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends.
The new methods developed in this paper yield measures that call into question a fact at the center of debates involving health, innovation, and fiscal policy. Discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press. Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a “warning sign” for the future of technological progress (Piper,, 2023).